[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue May 09, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed May 10 02:53:11 UTC 2017


HSU Shift Log for Tue May 09 2017 22:53:04 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: Unable to import satellite fixes from TAFB and SAB since  
09/18Z.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170510/ea503c8d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list